OpenStudy (anonymous):

solve by completing the square. x(to 2nd power) -6x-16=0

OpenStudy (anonymous):

The first thing to do is make it look like something you can complete the square for. Get the x^2 and the x part of the equation alone.

OpenStudy (cwtan):

Completing the square (x-3)^2-3^2-16=0

OpenStudy (cwtan):

(x-3)^2=25 x-3=\(\pm\)5

OpenStudy (anonymous):

thank you!!!